OpenStudy (anonymous):

Water issues from a hose pipe of internal diameter 1cm at a rate of 5 litres per minute. At what speed is the water flowing through the pipe??

OpenStudy (lgbasallote):

related rates?

OpenStudy (anonymous):

sorry

OpenStudy (lgbasallote):

so it is related rates

OpenStudy (lgbasallote):

^i meant calculus

OpenStudy (anonymous):

erm, maybe. but I have not started calculus yet.

OpenStudy (anonymous):

linear speed of water?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

divide the rate of water flow in volume by the area of the pipe

OpenStudy (anonymous):

why??

OpenStudy (anonymous):

area of pore of the pipe= pi*(1^2) sq cm= pi sq cm per minute water flows 5 litre= 5000 cc so, rate of flow of water = 5000/pi cm/minute= 1592 cm/minute (approx)

OpenStudy (anonymous):

volume= area*length so, volume/time= area* length/time=area* rate of water flow

OpenStudy (anonymous):

per minute or per second?

OpenStudy (anonymous):

per minute I believe

OpenStudy (anonymous):

sorry, the diameter is 1 cm ^^ so, the area is pi/4

OpenStudy (anonymous):

sorry I am really baffled

OpenStudy (anonymous):

me too.. after knowing the solution of ur text :)

OpenStudy (anonymous):

Actually the answer 106cm/s

OpenStudy (anonymous):

yeah, thats correct

OpenStudy (anonymous):

beg your pardon

OpenStudy (anonymous):

but I still do not fully understand.

OpenStudy (anonymous):

which part?

OpenStudy (anonymous):

area of pore of the pipe= pi*(1^2) sq cm= pi sq cm

OpenStudy (anonymous):

no, as 1 is the diameter, radius is 1/2 as area of the circle= pi r^2 so, area of the pore = pi/4

OpenStudy (anonymous):